Solve by using substitution or elimination:
x=3y+11
2x+5y=0

Answer by misscrt(37)

Solve by using substitution or elimination:
x=3y+11 (Eqn 1)
2x+5y=0 (Eqn 2)
_______________________________________________________
Solve by substitution(substitute expression for x found in Eqn 1 into Eqn 2:
2(3y+11)+5y=0
Use distributive property to clear the parentheses with multiplication:
2*3y + 2*11 + 5y = 0
6y+22+5y=0
Combine like terms:
6y+5y +22=0
11y+22=0
Subtract 22 from both sides to isolate the y-term:
11y+22-22=0-22
11y = -22
Divide both sides by 11 to get y:
11y/11=-22/11
y = -2
Find x value by plugging in y into Eqn 1:
x = 3(-2) + 11
x = -6 + 11
x = 5
Hope that helps.
Greetings,
misscrt